# Correlations Are Ruining Your Gradient Descent

### v2

@@ Appendix E Architectures and Hyperparameters
Dense Architecture Input 32$\times$32$\times$3
@@
Convolutional Architecture (CIFAR10/100) Input 32$\times$32$\times$3
Conv 3$\times$3$\times$32, (1, 0)
Conv 3$\times$3$\times$32, (1, 0)
MaxPool 2$\times$2 (2, 0)
Conv 3$\times$3$\times$64, (1, 0)
Conv 3$\times$3$\times$64, (1, 0)
MaxPool 2$\times$2 (2, 0)
@@
Convolutional Architecture (Tiny ImageNet) Input 56$\times$56$\times$3
Conv 3$\times$3$\times$32, (1, 0)
Conv 3$\times$3$\times$32, (1, 0)
MaxPool 2$\times$2 (2, 0)
Conv 3$\times$3$\times$64, (1, 0)
Conv 3$\times$3$\times$64, (1, 0)
MaxPool 2$\times$2 (2, 0)
@@
The CIFAR10(100) dataset is composed of a total training set of 50,000 samples of images (split into either 10 or 100 classes) and 10,000 test images. The TinyImageNet dataset is a sub-sampling of the ILSVRC (ImageNet) dataset, composed of 100,000 training images and 10,000 test images of 200 unique classes. Images from the ILSVRC dataset have been downsampled to 64$\times$64$\times$3 pixels and during our training and testing we further crop these images to 56$\times$56$\times$3 pixels. Cropping is carried out randomly during training and in a center-crop for testing.
@@
The training hyper-parameters were largely fixed across simulations, with a fixed (mini)batch size of 256, and Adam optimiser parameters of $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=1e-8$. Aside from these parameters, the learning rates for each simulation curve in Figure 4 were individually optimised. A single run of each simulation (curve), with a 10,000 sample validation set extracted from the training data, was run across a range of learning rates for both the forward and decorrelation optimisation independently. Learning rates were tested from the set $[1e-2,1e-3,1e-4,1e-5,1e-6,1e-7]$ for the forward learning rates, with and without decorrelation learning (again with learning rates tested from this set). Best parameters were selected for each simulation and thereafter the final results plots created based upon the full training and test sets and with five randomly seeded network models for each curve (see min and max performance as the envelopes shown in Figure 4.

### v3

@@ Appendix F Architectures and Hyperparameters
Dense Architecture Input $\times32$$\times$3
@@
Convolutional Architecture (CIFAR10/100) Input $\times32$$\times$3
Conv $\times3$$\times$32, (1, 0)
Conv $\times3$$\times$32, (1, 0)
MaxPool $\times2$ (2, 0)
Conv $\times3$$\times$64, (1, 0)
Conv $\times3$$\times$64, (1, 0)
MaxPool $\times2$ (2, 0)
@@
Convolutional Architecture (Tiny ImageNet) Input $\times56$$\times$3
Conv $\times3$$\times$32, (1, 0)
Conv $\times3$$\times$32, (1, 0)
MaxPool $\times2$ (2, 0)
Conv $\times3$$\times$64, (1, 0)
Conv $\times3$$\times$64, (1, 0)
MaxPool $\times2$ (2, 0)
@@
The CIFAR10(100) dataset is composed of a total training set of 50,000 samples of images (split into either 10 or 100 classes) and 10,000 test images. The TinyImageNet dataset is a sub-sampling of the ILSVRC (ImageNet) dataset, composed of 100,000 training images and 10,000 test images of 200 unique classes. Images from the ILSVRC dataset have been downsampled to $\times64$$\times$3 pixels and during our training and testing we further crop these images to $\times56$$\times$3 pixels. Cropping is carried out randomly during training and in a center-crop for testing.
@@
The training hyper-parameters were largely fixed across simulations, with a fixed (mini)batch size of 256, and Adam optimiser parameters of $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=1e-8$. Aside from these parameters, the learning rates for each simulation curve in Figure 4 were individually optimised. A single run of each simulation (curve), with a 10,000 sample validation set extracted from the training data, was run across a range of learning rates for both the forward and decorrelation optimisation independently. Learning rates were tested from the set [1e2,1e3,1e4,1e5,1e6,1e7][1e-2,1e-3,1e-4,1e-5,1e-6,1e-7] for the forward learning rates, with and without decorrelation learning (again with learning rates tested from this set). Best parameters were selected for each simulation and thereafter the final results plots created based upon the full training and test sets and with five randomly seeded network models for each curve (see min and max performance as the envelopes shown in Figure 4.